OpenStudy (anonymous):

Please help, been stuck on this and need a good grade, super confused. Please When solving a radical equation, Amari and Justin came to two different conclusions. Amari found a solution, while Justin's solution did not work in the equation. Create and justify two situations: one situation where Amari is correct and a separate situation where Justin is correct.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Not sure what they are expecting. Is this about extraneous solutions when solving radical equations?

OpenStudy (anonymous):

I think so, I think they're saying that Amari had a solution and that Justin didn't. so we would need extraneous and non-extraneous

OpenStudy (ranga):

Solve: \[\sqrt{x-2} = 5\] Square both sides: x - 2 = 25 x = 25 + 2 = 27 Put x = 27 back in the original equation and see if the solution is correct: sqrt(27-2) = sqrt(25) = 5 which is the same as the right hand side and therefore the solution is correct. Next take the radical equation: ...

OpenStudy (anonymous):

So that could be Amari?

OpenStudy (ranga):

Yes. The solution worked and that could be Amari.

OpenStudy (ranga):

Solve: \[\sqrt{x-1} = x- 7\] Square both sides: x - 1 = (x-7)^2 x - 1 = x^2 - 14x + 49 x^2 - 14x - x + 49 + 1 = 0 x^2 - 15x + 50 = 0 (x-10)(x-5) = 0 x = 5 or x = 10 Put each x back in the original equation and see if they satisfy: x = 5 LHS: sqrt(5-1) = 2 RHS: 5 - 7 = -2 LHS is not equal to RHS. So x = 5 is an extraneous solution introduced by squaring. Try x = 10. LHS: sqrt(10-1) = sqrt(9) = 3 RHS: 10 - 7 = 3 LHS = RHS. x = 10 is a solution. So one of Justin's solution, x = 5, did not work.

OpenStudy (anonymous):

thak you so much, I appreciate it I've been having a hard time!

OpenStudy (ranga):

You are welcome.